OpenStudy (enricoanders):

Two cyclists, 45 miles apart, start riding toward each other at the same time. One cycles 2 times as fast as the other. If they meet 3 hours later, what is the speed (in mi/h) of the faster cyclist? a. Write an equation using the information as it is given above that can be solved to answer this problem. Use the variable r to represent the speed of the slower cyclist.

OpenStudy (enricoanders):

@DanJS

OpenStudy (danjs):

If you say 'r' is the speed of the slower of the two, One travels at r, and the other at 2r. They start 45 mi apart and go towards the meeting point. The distance is being covered at the sum of their speeds , (r + 2r)

OpenStudy (enricoanders):

so its just 45=(r+2r)

OpenStudy (enricoanders):

i learned how to do this with chart

OpenStudy (danjs):

Remember that the distance traveled is the rate of travel times travel time. D = r * t

OpenStudy (danjs):

D = (r + 2r) *t D = (3r) *t They say they start at d=45 miles, and complete it at t=3 hr. 45 = 3*r * 3 45=9r r = 5 mi/hr

OpenStudy (enricoanders):

yeah that what i learned o the chart so its almost the same

OpenStudy (danjs):

the slower one goes 5 mi/hr, the faster one is 2r = 10 mi/hr